Can you please help me with this word problem?
The gravitational attraction between two masses varies inversely as the square of the distance between them. The force of attraction is 4 lbs. when the masses are 3 ft apart, what is the attraction when the masses are 6 ft. apart?
Not sure how to even set up, algebra is hard for me!

Answer by nyc_function(2741)

Let g = gravitational attraction force in terms of pounds between two masses.
Your equation is: g = k/d^2, where k is the constant of proportionality and d is our distance squared.
We know that g = 4.
We also know that 3 = d.
We first need to find k.
4 = k/(3)^2
4 = k/9
4*9 = k
36 = k
Now that we know k, we need to find g.
g = 36/(6)^2
g = 36/36
g = 1 pound.